OpenStudy (bookworm14):

will you Check my answer please? It is about the quadratic formula! (you will have to give me a min to draw it.)

OpenStudy (bookworm14):

I got \[9+5\sqrt{29} ~~and~~ 9-5\sqrt{29}\] Is that the final answer or can it be simplified to \[\sqrt{29} ~~and~~ -\sqrt{29}\]

OpenStudy (bookworm14):

the problem is \[7r^2-9r-23=0\]

OpenStudy (phi):

it looks like you forgot to divide by (2a) which is 2*7 = 14 in this problem \[ x = \frac{ 9 \pm 5\sqrt{29}}{14} \]

OpenStudy (bookworm14):

I was just about to type that into my questionxD I forgot to put it on here my bad haha

OpenStudy (bookworm14):

I have it all worked out on paper, I just did not enter it right on here on accident xD

OpenStudy (phi):

you can use a calculator and change both expressions to decimals and no, neither root can be simplified to sqr(29) (which you can also change to a decimal if you want to compare it to the roots)

OpenStudy (bookworm14):

i didn't know if the 9+5 (14) could be divided by the bottom 14 or not

OpenStudy (phi):

the approximate value of the roots are -1.2804 and 2.5661 sqr(29) is about 5.385

OpenStudy (bookworm14):

oh okay! thank you Phi !

OpenStudy (phi):

**i didn't know if the 9+5 (14) could be divided by the bottom 14 or not** do you mean 9 + 5 * sqr(29) ? yes , you can divide that by 14: you write a line under it, then 14 under the whole thing \[ \frac{9+5\sqrt{29}}{14} \] if you like, practice changing that into a decimal using a calculator.

OpenStudy (bookworm14):

so if I add (on the top) 5+19 that would become 14 sqrt 20 / 14 can the top 14 and the bottom 14 cancel each other out ? @phi

OpenStudy (phi):

you don't have 5+19 up top, you have \[ 9 + 5 \sqrt{29}\] and that is divided by 14 : \( \frac{9 + 5 \sqrt{29}}{14}\) It really can't be simplified (except to write an *approximate* decimal value) The reason why is this: First, remember sqrt(29) is a number (a bit larger than 5) so this problem is the same "form" as for example 1+2*3 all divided by 2: \( \frac{1+2\cdot 3}{2} \) you would not add 1+2 (I hope!). Order of operations says we do multiplies first. anyway, it would not turn into 3*3/2 = 9/2 it is (1+6)/2 = 7/2